i need help with some questions:
a triangular prism has the dimensions of HEIGHT 10M,LENGTH 12M, AND WIDTH 5M what is the volume.

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

volume of any prism is equal to the height times the area of the base.
the base can be a triangle or a rectangle or a polygon or a circle.
you need to find the area of the base first and then multiply that by the height.

if you have a triangular prism, then the base of the prism is a triangle.
a triangle has 3 sides.
length and width do not apply to a triangle.
you either need the 3 sides of the triangle or you need the width and height of the triangle.

if we assume you are given the width and height of the triangle rather than the width and length, then you would find the area of the triangle first and then multiply that by the height.

area of a triangle is equal to 1/2 times the base times the height.
assuming the base of your triangle is 12m and the height of your triangle is 5m, then the area of your triangle is 1/2 * 12 * 5 = 1/2 * 60 = 30

the volume of the triangular prism would be equal to the height of the prism times the area of the base which would be equal to 10 * 30 = 300 square cm.

here's a reference that might help you.
http://www.mathsisfun.com/geometry/prisms.html

here's a picture of your triangular prism.

as you can see, the length and width do not apply to a triangle.
length and width apply to a rectangle, but that's a different shape base.
the cross section in the reference is referring to the shape of the base.




Answer by ikleyn(53427)

.

A triangular prism does not have neither length, nor width.

Consider to visit special classes where they will teach you to write Math correctly.